# On total transitivity of graphs

### v1

@@ 5.3 Complexity Analysis
In the function Total_Transitive_Number(), we find the total transitive number or modified-total transitive number of a vertex $x$ based on the rooted total transitive number or rooted modified-total transitive number of its children. We assume the children are sorted according to their $l^{\prime}$-value, which is defined in Theorem 27. Since the for loop in lines $1-8$ of Total_Transitive_Number() runs for every child of $x$, this function takes $O(deg(x))$ time.
In the main algorithm Total_Transitivity(T), the vertex order mentioned in line $4$ can be found in linear time. Then, in a bottom-up approach, we calculate the rooted total transitive numbers or rooted modified-total transitive numbers of every vertex. For that, we are spending $O(deg(c_{i}))$ for every $c_{i}\in\sigma_{u}$. Note that we must pass the children of $c_{i}$ in a sorted order to Total_Transitive_Number(). But as discussed in [14] (an algorithm for finding the Grundy number of a tree), we do not need to sort all the children based on their rooted total transitive numbers or rooted modified-total transitive number; sorting the children whose rooted total transitive number or rooted modified-total transitive number is less than $deg(c_{i})$ is sufficient. We can argue that this can be done in $O(deg(c_{i}))$, as shown in [14]. Hence, the loop in lines $4-11$ takes linear time. To calculate the total transitive number for all vertices, we need $O(n^{2})$ times. Therefore, we have the following theorem:
@@ 6 Conclusion and future works
In this paper, we have studied the notion of total transitivity in graphs, which is a variation of transitivity. First, we have characterized split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Moreover, for split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we have shown some necessary conditions for $Tr_{t}(G)=p$. Furthermore, we have shown that the decision version of this problem is NP-complete for bipartite graphs. On the positive side, we have proved that this problem can be solved in polynomial time for trees.

### v2

@@ 6.3 Complexity Analysis
In the function Total_Transitive_Number(), we find the total transitive number or weak total transitive number of a vertex $x$ based on the rooted total transitive number or rooted weak total transitive number of its children. We assume the children are sorted according to their $l^{\prime}$-value, which is defined in Theorem 30. It is easy to see that the function Total_Transitive_Number() takes $O(deg(x))$ time for a vertex $x$.
In the main algorithm Total_Transitivity(T), the vertex order mentioned in line $4$ can be found in linear time. Then, in a bottom-up approach, we calculate the rooted total transitive numbers or rooted weak total transitive numbers of every vertex. For that, we are spending $O(deg(c_{i}))$ for every $c_{i}\in\sigma_{u}$. Note that we must pass the children of $c_{i}$ in a sorted order to Total_Transitive_Number(). But as discussed in [14] (an algorithm for finding the Grundy number of a tree), we do not need to sort all the children based on their rooted total transitive numbers or rooted weak total transitive number; sorting the children whose rooted total transitive number or rooted weak total transitive number is less than $deg(c_{i})$ is sufficient. We can argue that this can be done in $O(deg(c_{i}))$, as shown in [14]. Hence, the loop in lines $4-11$ takes linear time. To calculate the total transitive number for all vertices, we need $O(n^{2})$ times. Therefore, we have the following theorem:
@@ 7 Conclusion and future works
In this paper, we have studied the notion of total transitivity in graphs, which is a variation of transitivity. First, we have characterized split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Moreover, for split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we have shown some necessary conditions for $Tr_{t}(G)=p$. Furthermore, we have shown that the decision version of this problem is NP-complete for bipartite graphs. On the positive side, we have proved that this problem can be solved in linear time for bipartite chain graphs. Finally, we have designed a polynomial-time algorithm for trees.